# SKEL-CF: Coarse-to-Fine Biomechanical Skeleton and Surface Mesh Recovery

### v4

@@ Abstract
Parametric 3D human models such as SMPL have driven significant advances in human pose and shape estimation, yet their simplified kinematics limit biomechanical realism. The recently proposed SKEL model addresses this limitation by re-rigging SMPL with an anatomically accurate skeleton. However, estimating SKEL parameters directly remains challenging due to limited training data, perspective ambiguities, and the inherent complexity of human articulation. We introduce SKEL-CF, a coarse-to-fine framework for SKEL parameter estimation. SKEL-CF employs a transformer-based encoder–decoder architecture, where the encoder predicts coarse camera and SKEL parameters, and the decoder progressively refines them in successive layers. To ensure anatomically consistent supervision, we convert the existing SMPL-based dataset 4DHuman into a SKEL-aligned version, 4DHuman-SKEL, providing high-quality training data for SKEL estimation. In addition, to mitigate depth and scale ambiguities, we explicitly incorporate camera modeling into the SKEL-CF pipeline and demonstrate its importance across diverse viewpoints. Extensive experiments validate the effectiveness of the proposed design. On the challenging MOYO dataset, SKEL-CF achieves 85.0 MPJPE / 51.4 PA-MPJPE, significantly outperforming the previous SKEL-based state-of-the-art HSMR (104.5 / 79.6). These results establish SKEL-CF as a scalable and anatomically faithful framework for human motion analysis, facilitating the use of computer vision techniques in biomechanics-related analysis. Our implementation is available on the project page: https://pokerman8.github.io/SKEL-CF/.
Figure 1: Visualization of SKEL-CF on web images. Compared with SMPL-based state-of-the-art CameraHMR [27], our SKEL-based model produces more natural joint motions (see the side-view zoomed knee in blue). Compared with HSMR [39], the state-of-the-art SKEL-based method, SKEL-CF achieves more accurate skeleton and mesh reconstruction (see zoomed hand in orange).
@@ 1 Introduction
In recent years, 3D human pose and shape estimation has made remarkable progress, enabling a wide range of downstream applications [5, 8, 19, 37, 32]. However, adoption in biomechanics, a domain where such techniques could be particularly impactful, remains limited. This gap arises because existing parametric models, such as SMPL [23], SMPL-X [28], and GHUM [40], fall short of the stringent biomechanical requirements. Their simplified kinematics and loosely constrained axis–angle representations often compromise biomechanical realism, especially in complex articulations such as deep squats or yoga poses [39]. In these scenarios, the estimated parameters must correspond to anatomically accurate skeletons, respect joint limits, and ensure physically plausible motion with high precision. As illustrated in Figure 1 (a), even the state-of-the-art SMPL-based method CameraHMR [27] can produce anatomically implausible results, such as unnatural knee bending.
@@
To address these limitations, the recently proposed SKEL [15] model redefines the foundation of human representation by re-rigging SMPL with an anatomically accurate skeleton and realistic joint constraints. This makes it a promising step toward bridging computer vision and biomechanics, offering a representation that is both visually coherent and biomechanically faithful. HSMR [39] is the first transformer-based method to reconstruct a human body in the SKEL space from a single image. It converts the predicted SMPL ground truth from existing datasets into SKEL pseudo ground truth for supervised training. However, estimating SKEL from a single image remains challenging, as it requires inferring 3D structure from inherently limited 2D observations. Consequently, the predictions often lack accuracy (see Figure 1 (b)), and empirical metrics such as MPJPE remain inferior to SMPL-based models.
@@
In this work, we take a step forward in advancing SKEL-based human mesh recovery. We observe that estimating SKEL parameters in a coarse-to-fine manner significantly improves performance. Specifically, our transformer encoder first predicts coarse camera and SKEL parameters, while the decoder refines these estimates progressively, layer by layer, leading to more accurate and stable results. To address challenges arising from diverse camera viewpoints, we incorporate camera intrinsic estimation following CameraHMR [27], which enhances the robustness of the reconstructed mesh across varying perspectives. Moreover, by leveraging the high-quality SMPL estimations provided by the recent 4DHuman dataset, we generate large-scale, high-fidelity SKEL annotations to form the 4DHuman-SKEL dataset, enabling more reliable training for SKEL estimation.
@@
Empirically, our proposed SKEL-CF achieves substantial performance gains across diverse benchmarks. On the challenging MOYO dataset [33], SKEL-CF reaches 85.0 MPJPE and 51.4 PA-MPJPE, significantly outperforming the previous SKEL-based state-of-the-art HSMR [39], which reports 104.5 and 79.6, respectively. Furthermore, SKEL-CF attains comparable or superior quantitative performance to leading SMPL-based approaches such as CameraHMR [27] across multiple datasets, including 3DPW [35], EMDB [14], and MOYO [33], while producing biomechanically faithful and visually consistent mesh reconstructions (See igure 1 (c)) These results position SKEL-CF as a robust and anatomically realistic solution for practical human mesh recovery in real-world scenarios. We also conduct ablation studies to validate the contribution of each proposed component in our framework.
@@
  • We propose SKEL-CF, a coarse-to-fine transformer framework that significantly advances SKEL-based human mesh recovery. By explicitly modeling the camera and refining SKEL parameters layer by layer, SKEL-CF achieves state-of-the-art performance across multiple benchmarks while maintaining biomechanical fidelity.
  • We demonstrate that estimating SKEL parameters in a coarse-to-fine and iterative refinement manner leads to substantial performance gains. The decoder progressively improves predictions across layers, effectively handling complex articulations and challenging poses.
  • We introduce 4DHuman-SKEL, a large-scale, high-quality dataset with anatomically consistent SKEL annotations converted from 4DHuman [6]. This dataset provides reliable supervision and unified skeleton–mesh alignment, enabling robust SKEL-based model training.
Our implementation is available on the project page: https://pokerman8.github.io/SKEL-CF/.
Figure 2: Overview of the proposed SKEL-CF. Our method estimates SKEL parameters from a single image using an encoder–decoder architecture that performs coarse-to-fine estimation. The encoder produces initial predictions of the camera extrinsics $\boldsymbol{\pi}$, shape parameters $\boldsymbol{\beta}$, and pose parameters $\boldsymbol{\theta}$. The decoder then refines these predictions progressively across layers. In addition, we adopt the pretrained camera model from CameraHMR [27] to estimate the camera intrinsics, and keep its parameters frozen throughout training.
@@ 2 Related Work
Parametric human body models are widely used for 3D human shape and pose estimation, as they offer a structured low-dimensional representation of human geometry. SMPL [23] represents the human mesh with fixed topology, using low-dimensional shape and pose parameters with linear blend skinning and pose-dependent corrective shapes for realistic articulation. SMPL-H [29] augments this framework with articulated hands, while SMPL-X [28] further integrates the FLAME [20] facial model, yielding a unified representation of body, hands, and expressive face. While the SMPL family provides a compact and effective representation of the human body surface, it employs a simplified kinematic structure that deviates from the anatomical skeletal system of real humans, allowing anatomically implausible poses (e.g., the knee can exhibit unrealistically free rotation, see Figure 1 (b)). SKEL [15] introduces anatomically accurate joint definitions and a bone hierarchy embedded within a parametric human body model, overcoming limitations of simplified kinematic structures in models such as SMPL. It can produce a skeleton mesh aligned with the body surface that is suitable for applications requiring anatomically valid joint locations and realistic bone articulations, such as biomechanical analysis or as a basis for further physics-based simulation.
@@
The lack of 3D annotations for in-the-wild images remains a major obstacle in Human mesh recovery. SMPLify [2] mitigates this by iteratively optimizing SMPL [23] parameters from 2D keypoints, yielding accurate results but at high computational cost. HMR [11] introduces the first end-to-end framework to predict 3D pose and shape from in-the-wild images using only 2D supervision, where adversarial learning with mocap data provides strong pose and shape priors to compensate for the missing 3D information. 4DHuman [6] constructs a large-scale dataset by fitting approximately 3 million images with ProHMR [18], generating pseudo-ground-truth (p-GT) 3D annotations alongside 2D keypoints, and further upgrades the network architecture from CNN to Vision Transformer (ViT). However, TokenHMR [4] identifies a mismatch between 3D and 2D keypoints, where enforcing 2D keypoint alignment can degrade 3D evaluation metrics, partly due to the use of fixed camera intrinsics and p-GT annotations in the 4DHuman [6] dataset. To address this, TokenHMR [4] introduces the TALS loss, which down-weights the 2D keypoint loss when the L2 distance between predicted and ground-truth keypoints falls within a predefined threshold, thus preventing overfitting to noisy p-GT annotations. In addition, it leverages a quantized token dictionary, constructed by pre-training a Vector Quantized-VAE (VQ-VAE) [34] on motion capture datasets [25, 33], to mitigate 3D pose ambiguity. CameraHMR [27] improves the 4DHuman [6] dataset through CamSMPLify, which refines body representations using dense keypoints to overcome the average body limitation. To tackle the fixed-camera constraint, it introduces HumanFOV [27], transforming weak-perspective projection into a fully perspective projection.
@@
The first large-scale 3D pseudo-ground-truth (p-GT) dataset was introduced by 4DHuman [6], which extends to unlabeled datasets such as InstaVariety [13], AVA [7], and AI Challenger [31] by generating p-GT annotations. For each image, an off-the-shelf detector [3] and a body keypoint estimator are applied to obtain bounding boxes and 2D keypoints. A SMPL [23] mesh is then fitted to these keypoints using ProHMR [18], producing pseudo-ground-truth SMPL parameters. Despite its scale and utility, the 4DHuman dataset still contains inaccuracies and artifacts. To that end, CameraHMR [27] refines the dataset using its proposed CamSMPLify method, alongside additional techniques to enhance annotation quality. While large-scale datasets of SMPL parameters have been constructed, there remains a need to develop datasets for SKEL [15] parameters. HSMR [39] addresses this by first converting the 4DHuman SMPL parameters to SKEL parameters using SKEL fitting [15]. To mitigate the inaccuracies in 4DHuman annotations, we further apply SKEL fitting to the refined SMPL annotations provided by CameraHMR.
@@ 3 Method
In this section, we introduce our approach, SKEL-CF, whose overall pipeline is illustrated in Figure 2. SKEL-CF aims to estimate SKEL parameters from a single image using an encoder–decoder architecture that performs coarse-to-fine estimation. The encoder produces initial predictions of the camera extrinsics $\boldsymbol{\pi_{0}}$, shape parameters $\boldsymbol{\beta_{0}}$, and pose parameters $\boldsymbol{\theta_{0}}$, while the decoder progressively refines these predictions across layers. To enhance robustness to camera variations, we adopt the pretrained camera model from CameraHMR [27] to estimate the camera intrinsics, and keep its parameters frozen throughout training.
This section is organized as follows. We first introduce the SKEL model and the 4DHuman-SKEL dataset that we construct to supervise SKEL learning in Section 3.1. We then describe the details of SKEL-CF and its implementation in Section 3.2.
@@ 3.1 Preliminaries
The SKEL model [15] is a parametric representation of the human body that enables the joint reconstruction of both the skin surface and the skeletal mesh. Given the pose parameters $q\in\mathbb{R}^{46}$ and the shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ (corresponding to the top 10 shape components), SKEL generates the human mesh $M\in\mathbb{R}^{6890\times 3}$ and the skeleton mesh $S_{k}\in\mathbb{R}^{24752\times 3}$. Unlike SMPL, which defines pose parameters $\theta\in\mathbb{R}^{72}$ using unrestricted three-degree-of-freedom ball joints, SKEL constrains the motion of each joint, reducing the parameter space from 72 to 46. This formulation ensures anatomically consistent and biomechanically valid poses [15, 39]. Consequently, SKEL produces physically plausible human motion and is particularly suitable for downstream applications such as motion analysis, rehabilitation, and human–robot interaction.
4DHuman-SKEL: a high-quality dataset for SKEL training.
HSMR [39] initially generated pseudo ground-truth (p-GT) SKEL parameters by converting the SMPL parameters into the SKEL parameter space. While this conversion provided a practical starting point, its quality was limited by the low-resolution images and noisy annotations in the original 4DHuman [6] dataset, which aggregates samples from Human3.6M [9], MPI-INF-3DHP [26], COCO [22], MPII [1], AI Challenger [38], AVA [7] and InstaVariety [12]. With the recent release of the refined 4DHuman SMPL annotations [6] from CameraHMR [27], we revisit this process to obtain higher-quality SKEL p-GT annotations. Consequently, we exclude AVA, as CameraHMR does not provide refined SMPL annotations for this subset, and denote the resulting dataset as 4DHuman-SKEL.
Specifically, we reconstruct the human mesh from the provided SMPL [23] parameters using the neutral SMPL model. Our overall optimization pipeline largely follows the SKEL [15] fitting protocol (i.e., starting from an initial set of SKEL parameters and performing iterative optimization with a hierarchical refinement schedule), but we introduce a small yet important adaptation to better align with the refined 4DHuman dataset [6]. Different from SKEL, since both models share identical global orientation and translation in our setting, we directly convert the SMPL global orientation from axis–angle to Euler angles to comply with the SKEL convention and fix these parameters during optimization. We observe that further optimizing or re-learning global orientation and translation does not improve alignment quality on the refined dataset. This process takes approximately 58 hours to process all 3M images on a single RTX 3090 GPU.

### v5

@@ Abstract
Parametric 3D human models such as SMPL have driven significant advances in human pose and shape estimation, yet their simplified kinematics limit biomechanical realism. The recently proposed SKEL model addresses this limitation by re-rigging SMPL with an anatomically accurate skeleton. However, estimating SKEL parameters directly remains challenging due to limited training data, perspective ambiguities, and the inherent complexity of human articulation. In this work, we propose SKEL-CF, a new framework for estimating SKEL parameters. SKEL-CF adopts a standard transformer-based encoder–decoder architecture. The encoder first produces coarse predictions of the camera extrinsics and SKEL parameters. The decoder then iteratively refines these predictions across multiple layers, with explicit and auxiliary supervision applied at each layer. To provide anatomically consistent training data, we convert the existing SMPL-based dataset into a SKEL-aligned version, called HMR-SKEL. This new dataset offers high-quality supervision for SKEL estimation. In addition, to reduce depth and scale ambiguity, we explicitly incorporate camera intrinsic estimation into the SKEL-CF pipeline and show that it is important for accurate reconstruction. Extensive experiments validate the effectiveness of the proposed design. On the challenging MOYO dataset, SKEL-CF achieves 85.0 MPJPE / 51.4 PA-MPJPE, significantly outperforming the previous SKEL-based state-of-the-art HSMR (104.5 / 79.6). These results establish SKEL-CF as a scalable and anatomically faithful framework for human motion analysis, facilitating the use of computer vision techniques in biomechanics-related analysis. Our implementation is available on the project page: https://pokerman8.github.io/SKEL-CF/.
@@ 1 Introduction
Figure 1: Visualization of SKEL-CF on web images. Compared with SMPL-based state-of-the-art CameraHMR [patel2024camerahmr], our SKEL-based model produces more natural joint motions (see the side-view zoomed knee in blue). Compared with HSMR [xia2025hsmr], the state-of-the-art SKEL-based method, SKEL-CF achieves more accurate skeleton and mesh reconstruction (see zoomed hand in orange).
@@
In recent years, 3D human pose and shape estimation has made remarkable progress, enabling a wide range of downstream applications [fu2024humanplus, he2024human2humannoid, li2024ohumanoidrobot, weng2022humannerf, tevet2022human]. However, adoption in biomechanics, a domain where such techniques could be particularly impactful, remains limited. This gap arises because existing parametric models, such as SMPL [loper2015smpl], SMPL-X [smpl-x], and GHUM [xu2020ghum], fall short of the stringent biomechanical requirements. Their simplified kinematics and loosely constrained axis–angle representations often compromise biomechanical realism, especially in complex articulations such as deep squats or yoga poses [xia2025hsmr]. In these scenarios, the estimated parameters must correspond to anatomically accurate skeletons, respect joint limits, and ensure physically plausible motion with high precision. As illustrated in Figure 1 (a), even the state-of-the-art SMPL-based method CameraHMR [patel2024camerahmr] can produce anatomically implausible results, such as unnatural knee bending.
@@
To address these limitations, the recently proposed SKEL [SKEL] model redefines the foundation of human representation by re-rigging SMPL with an anatomically accurate skeleton and realistic joint constraints. This makes it a promising step toward bridging computer vision and biomechanics, offering a representation that is both visually coherent and biomechanically faithful. HSMR [xia2025hsmr] is the first transformer-based method to reconstruct a human body in the SKEL space from a single image. It converts the predicted SMPL ground truth from existing datasets into SKEL pseudo ground truth for supervised training. However, estimating SKEL from a single image remains challenging because it requires recovering a more constrained 3D structure than SMPL under the inherent depth ambiguity of monocular input. Consequently, the predictions often lack accuracy (see Figure 1 (b)), and empirical metrics such as MPJPE remain inferior to SMPL-based models.
@@
In this work, we advance SKEL-based human mesh recovery by proposing SKEL-CF, a new framework for estimating SKEL parameters. Similar to HSMR [xia2025hsmr], SKEL-CF follows a standard transformer-based encoder–decoder architecture. However, unlike HSMR, which applies supervision only at the final layer, we adopt a coarse-to-fine estimation strategy. The encoder first predicts coarse camera extrinsics and SKEL parameters. The decoder then refines these predictions. Even a single refinement step yields noticeable performance gains. Moreover, by iteratively refining the predictions layer by layer in the decoder, with supervision applied at each stage, we further improve the performance. To handle diverse camera viewpoints, we incorporate camera intrinsic estimation following CameraHMR [patel2024camerahmr]. This improves robustness across different perspectives and reduces depth and scale ambiguity. Finally, leveraging the high-quality SMPL estimations from CameraHMR [patel2024camerahmr], we generate large-scale and high-fidelity SKEL annotations to construct HMR-SKEL. This dataset provides reliable supervision and enables more effective training for SKEL estimation.
@@
Empirically, our proposed SKEL-CF achieves substantial performance gains across diverse benchmarks. On the challenging MOYO dataset [moyo], SKEL-CF reaches 85.0 MPJPE and 51.4 PA-MPJPE, significantly outperforming the previous SKEL-based state-of-the-art HSMR [xia2025hsmr], which reports 104.5 and 79.6, respectively. Furthermore, SKEL-CF attains comparable or superior quantitative performance to leading SMPL-based approaches such as CameraHMR [patel2024camerahmr] across multiple datasets, including 3DPW [3dpw], EMDB [kaufmann2023emdb], and MOYO [moyo], while producing biomechanically faithful and visually consistent mesh reconstructions (See Figure 1 (c)) These results position SKEL-CF as a robust and anatomically realistic solution for practical human mesh recovery in real-world scenarios. We also conduct ablation studies to validate the contribution of each proposed component in our framework.
@@
  • We propose SKEL-CF, a new framework that significantly advances SKEL-based human mesh recovery. SKEL-CF achieves state-of-the-art performance across multiple benchmarks while maintaining biomechanical fidelity.
  • We show that estimating SKEL parameters in a coarse-to-fine manner, together with explicitly modeling camera intrinsics, leads to significant performance improvements.
  • We introduce HMR-SKEL, a large-scale, high-quality dataset with anatomically consistent SKEL annotations converted from smpl-based dataset [patel2024camerahmr]. This dataset provides reliable supervision and unified skeleton–mesh alignment, enabling robust SKEL-based model training.
Our data and code will be released upon publication.
@@ 2 Related Work
Parametric human body models are widely used for 3D human shape and pose estimation, as they offer a structured low-dimensional representation of human geometry. SMPL [loper2015smpl] represents the human mesh with fixed topology, using low-dimensional shape and pose parameters with linear blend skinning and pose-dependent corrective shapes for realistic articulation. SMPL-H [smpl-h] augments this framework with articulated hands, while SMPL-X [smpl-x] further integrates the FLAME [FLAME-HEAD] facial model, yielding a unified representation of body, hands, and expressive face. While the SMPL family provides a compact and effective representation of the human body surface, it employs a simplified kinematic structure that deviates from the anatomical skeletal system of real humans, allowing anatomically implausible poses (e.g., the knee can exhibit unrealistically free rotation, see Figure 1 (b)). SKEL [SKEL] introduces anatomically accurate joint definitions and a bone hierarchy embedded within a parametric human body model, overcoming limitations of simplified kinematic structures in models such as SMPL. By estimating parameters that are explicitly compatible with biomechanical skeletons, SKEL enforces realistic joint limits and produces physically plausible motion, leading to more accurate and reliable pose reconstruction.
@@
The lack of 3D annotations for in-the-wild images remains a major obstacle in Human mesh recovery. SMPLify [smplify] mitigates this by iteratively optimizing SMPL [loper2015smpl] parameters from 2D keypoints, yielding accurate results but at high computational cost. HMR [kanazawa2018hmr] introduces the first end-to-end framework to predict 3D pose and shape from in-the-wild images using only 2D supervision, where adversarial learning with mocap data provides strong pose and shape priors to compensate for the missing 3D information. 4DHuman [goel20234dhuman] constructs a large-scale dataset by fitting approximately 3 million images with ProHMR [prohmr], generating pseudo-ground-truth (p-GT) 3D annotations alongside 2D keypoints, and further upgrades the network architecture from CNN to Vision Transformer (ViT). However, TokenHMR [dwivedi2024tokenhmr] identifies a mismatch between 3D and 2D keypoints, where enforcing 2D keypoint alignment can degrade 3D evaluation metrics, partly due to the use of fixed camera intrinsics and p-GT annotations in the 4DHuman [goel20234dhuman] dataset. To address this, TokenHMR [dwivedi2024tokenhmr] introduces the TALS loss, which down-weights the 2D keypoint loss when the L2 distance between predicted and ground-truth keypoints falls within a predefined threshold, thus preventing overfitting to noisy p-GT annotations. In addition, it leverages a quantized token dictionary, constructed by pre-training a Vector Quantized-VAE (VQ-VAE) [van2017vqvae] on motion capture datasets [mahmood2019amass, moyo], to mitigate 3D pose ambiguity. CameraHMR [patel2024camerahmr] improves the 4DHuman [goel20234dhuman] dataset through CamSMPLify, which refines body representations using dense keypoints to overcome the average body limitation. To tackle the fixed-camera constraint, it introduces HumanFOV [patel2024camerahmr], transforming weak-perspective projection into a fully perspective projection.
@@
The first large-scale 3D pseudo-ground-truth (p-GT) dataset was introduced by 4DHuman [goel20234dhuman], which extends to unlabeled datasets such as InstaVariety [kanazawa2019insta], AVA [gu2018ava], and AI Challenger [sun2019aic] by generating p-GT annotations. For each image, an off-the-shelf detector [detr] and a body keypoint estimator are applied to obtain bounding boxes and 2D keypoints. A SMPL [loper2015smpl] mesh is then fitted to these keypoints using ProHMR [prohmr], producing pseudo-ground-truth SMPL parameters. Despite its scale and utility, the 4DHuman dataset still contains inaccuracies and artifacts. To that end, CameraHMR [patel2024camerahmr] refines the dataset using its proposed CamSMPLify method, alongside additional techniques to enhance annotation quality. While large-scale datasets of SMPL parameters have been constructed, there remains a need to develop datasets for SKEL [SKEL] parameters. HSMR [xia2025hsmr] addresses this by first converting the 4DHuman SMPL parameters to SKEL parameters using SKEL fitting [SKEL]. To mitigate the inaccuracies in 4DHuman annotations, we further apply SKEL fitting to the refined SMPL annotations provided by CameraHMR [patel2024camerahmr].
Figure 2: Overview of the proposed SKEL-CF. Our method estimates SKEL parameters from a single image using an encoder–decoder architecture that performs coarse-to-fine estimation. The encoder produces initial predictions of the camera extrinsics $\boldsymbol{\pi}$, shape parameters $\boldsymbol{\beta}$, and pose parameters $\boldsymbol{\theta}$. The decoder then refines these predictions iteratively across layers. In addition, we adopt the pretrained camera model from CameraHMR [patel2024camerahmr] to estimate the focal length, and keep its parameters frozen throughout training.
@@ 3 Method
In this section, we introduce our approach, SKEL-CF, whose overall pipeline is illustrated in Figure 2. SKEL-CF aims to estimate SKEL parameters from a single image using an encoder–decoder architecture that performs coarse-to-fine estimation. The encoder produces initial predictions of the camera extrinsics $\boldsymbol{\pi_{0}}$, shape parameters $\boldsymbol{\beta_{0}}$, and pose parameters $\boldsymbol{\theta_{0}}$, while the decoder iteratively refines these predictions across layers. To enhance robustness to camera variations, we adopt the pretrained camera intrinsic predictor from CameraHMR [patel2024camerahmr] to estimate the focal length, keeping its parameters frozen throughout training. This section is organized as follows. We first introduce the SKEL model and the HMR-SKEL dataset that we construct to supervise SKEL learning in Section 3.1. We then describe the details of SKEL-CF and its implementation in Section 3.2.
@@ 3.1 Preliminaries
The SKEL model [SKEL] is a parametric representation of the human body that enables the unified reconstruction of both the surface mesh and the skeleton mesh. Given the pose parameters $q\in\mathbb{R}^{46}$ and the shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ (corresponding to the top 10 shape components), SKEL generates the human mesh $M\in\mathbb{R}^{6890\times 3}$ and the skeleton mesh $S_{k}\in\mathbb{R}^{24752\times 3}$. Unlike SMPL, which represents pose parameters $\theta\in\mathbb{R}^{72}$ using unconstrained three-degree-of-freedom ball joints, SKEL constrains the motion of each joint. This reduces the pose parameter space from 72 to 46. For example, SKEL models the elbow as a hinge joint by limiting its degrees of freedom. This formulation ensures anatomically consistent and biomechanically valid poses [SKEL, xia2025hsmr]. Consequently, SKEL produces physically plausible human motion and is particularly suitable for downstream applications such as motion analysis, rehabilitation, and human–robot interaction.
HMR-SKEL: a high-quality dataset for SKEL training.
HSMR [xia2025hsmr] generated pseudo ground-truth (p-GT) SKEL parameters by converting SMPL parameters into the SKEL space. While this provided a practical starting point, its quality was limited by low-resolution images and noisy SMPL annotations in the original dataset. With the release of refined SMPL annotations from CameraHMR [patel2024camerahmr], we revisit this conversion process to obtain higher-quality SKEL pseudo ground truth. Following the HSMR setup, which is based on HMR2.0 [goel20234dhuman], we use data from Human3.6M [human36m], MPI-INF-3DHP [mpi-inf], COCO [coco], MPII [mpii], AI Challenger [aic], and InstaVariety [insta]. We exclude AVA because CameraHMR does not provide refined SMPL annotations for this subset. We denote the resulting dataset as HMR-SKEL.
Our optimization pipeline follows the fitting protocol of SKEL [SKEL]. We first reconstruct a reference human mesh from the high-quality annotations provided by CameraHMR [patel2024camerahmr], which is used as the target SMPL mesh. Given an initial set of SKEL parameters, we treat them as learnable variables and feed them into the SKEL model to generate a SKEL mesh. We then compute the alignment loss between the generated SKEL mesh and the target SMPL mesh, and use this loss to iteratively update the SKEL parameters. To improve optimization stability, we perform the fitting in a hierarchical manner: we first optimize the lower-body parameters, then the upper-body parameters, and finally the full-body parameters. Moreover, since both SMPL and SKEL share the same global orientation and translation terms, we keep these variables fixed during optimization; empirically, allowing them to vary leads to worse alignment. The full process takes about 58 hours to process 3M images on a single RTX 3090 GPU.